OpenStudy (anonymous):

What is the lim(x^2-25)/(x-5) as x approaches 5

Directrix (directrix):

10

OpenStudy (anonymous):

can you explain it im having trouble understanding limits

OpenStudy (accessdenied):

Factor the numerator into (x+5)(x-5). The like expressions cancel out, leaving x+5. Then, just plug and play.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 5} \frac{\cancel{(x-5)}(x+5)}{\cancel{(x-5)}}\]

Directrix (directrix):

m(x^2-25)/(x-5) as x approaches 5 = lim (x+5)(x-5) / (x-5) = lim (x+5) as x approaches 5 = lim ( (5) +5) = lim (10) as x approaches 5 = 10. Check out the graph at this link: http://www.wolframalpha.com/input/?i=y%3D+%28x%5E2-25%29%2F%28x-5%29